OpenStudy (unklerhaukus):

\[\int ax^n\text dx\] let\[ax^n=u\]\[x=\left(\frac{u}{a}\right)^{1/n}=\frac{u^{1/n}}{a^{1/n}}\]\[\text dx=\frac{u^{1/n-1}}{na^{1/n}}\text du\] \[\int ax^n\text dx\longrightarrow\frac{1}{na^{1/n}}\int u\cdot{u^{1/n-1}}\text du\] \[=\frac{1}{na^{1/n}}\int {u^{1/n}}\text du\]\[=\frac{1}{na^{1/n}}\frac{u^{1/n+1}}{1/n+1}+c\]\[={}\frac{a^{1+n}x^{1+n}}{a^{1/n}+na^{1/n}}+c\]\[=\frac{ax^{1+n}}{1+n}+c\]

OpenStudy (unklerhaukus):

what can i do with this

OpenStudy (anonymous):

integrate stuff ? for example \[\int\limits^{}3x^{2}=x^{3}+C\]

OpenStudy (anonymous):

Why would you integrate \(ax^n\) like that... Just use the linearity of integrals and the "power rule".

OpenStudy (unklerhaukus):

i know right this is a one step problem , i realized this when i got the the end, is there any use you can think for all this latex?

mathslover (mathslover):

your question is to integrate ax^n dx ?

OpenStudy (mathmate):

Perhaps in recognizing terms in summing series.

OpenStudy (unklerhaukus):

pardon @mathmate ?

OpenStudy (mathmate):

What I mean is, when we sum series, sometimes it is helpful to recognize terms in different forms. If each term is the integral / derivative of the terms of an expansion of a particular function, then the sum of the series is the integral/derivative of that function.

OpenStudy (anonymous):

Is Suggestion which is a little vague accepted ?

OpenStudy (anonymous):

@UnkleRhaukus ?

OpenStudy (unklerhaukus):

?

OpenStudy (anonymous):

You are applying an holomorphic function to the variable - lets extend it to complex domain:

OpenStudy (anonymous):

\[z^{\frac{1}{n}} = e^{\frac{1}{n}\log z} \]

OpenStudy (unklerhaukus):

ok , not too fast now

OpenStudy (anonymous):

So NOW you are integrating on a differently-shaped domain in the complex plane - therefore you can do a lot of integrals differently - not this "naked" function alone but, for example , multiplied by some other function which looks simpler in the new representation

OpenStudy (anonymous):

These geometric- holomorphic transforms are widely used in some areas of science and engineering

OpenStudy (anonymous):

Look up antenna theory, Greens functions [ in many areas - wave propagation, solid state, QM, Field theory etc]

OpenStudy (anonymous):

Mr champion ?

OpenStudy (unklerhaukus):

go on

OpenStudy (anonymous):

That all seems consistent, but quite unnecessary.

OpenStudy (anonymous):

I appologised from the start -" Suggestion which is a little vague ". I saw quite MASSIVE use of such domain transformation in some calculation in electromagnetism, and Wave propagation, and Green's functions computing generally. But it was some years ago, and I don't recall more precise area. Look up "geometric transformations of the complex plane" and Green function/antenna theory/etc

OpenStudy (anonymous):

@UnkleRhaukus ghmm .....

OpenStudy (anonymous):

ghhm Ghhm GHMM-MM-MM !

OpenStudy (unklerhaukus):

complexification

OpenStudy (anonymous):

what about some (purely symbolic) gratification ...

OpenStudy (unklerhaukus):

im so confused

OpenStudy (anonymous):

*is also confused* What is going on here? \[\int\limits ax^n dx = a \int\limits x^n dx = a \cdot \frac{x^{n+1}}{n+1} +c.\] by power rule, so what's the issue?

OpenStudy (anonymous):

He showed how to compute it by a change of variables and asked whether it can be interesting/useful . I gave him the interpretation where it is used and he (good FREE lesson on human nature...) did NOT thank me...

OpenStudy (anonymous):

Another good free lesson on human nature is how often you outright ask for medals and fans . . . ;-)